# Diffusion Models Are Promising for Ab Initio Structure Solutions from Nanocrystalline Powder Diffraction Data

### v1

@@ Results
Figure 9 shows $R_{wp}^{2}$ values for these candidates before and after the Rietveld refinement step. The Rietveld refinement works well for the 100 Å test cases, which have sharper Bragg peaks. Except for four cases, the refinement brought $R_{wp}^{2}$ below 20%. There may still be some issues with the structure in these cases, but PXRDnet is clearly returning structures that are quite similar to the true one and possibly close enough to yield a correct structure with some human input in each case.
We note that the average unrefined $R_{wp}^{2}$ of this sampling of materials were actually worse than the unrefined $R_{wp}^{2}$ of the materials in the overall testing set ($0.526$ vs. $0.377$ for 10 Å crystals; $0.478$ vs. $0.364$ for 100 Å crystals). The subset of targets we chose for refinement are not best-case or even average-case, and we might expect Rietveld refinements on wider ranges of targets would perform better than this sampling on average.
@@
The Rietveld refinements were much less successful for the 10 Å nanoparticles, which have highly distorted diffraction patterns. In many cases the refinement resulted in a worse $R_{wp}^{2}$, indicating that the refinements were not converged and not stable.
@@
This aspect of the work can be developed in the future, with better understanding of how to refine structural models in a more automated way, either against the PDF or against powder data, to obtain a higher structure solution rate.
@@ Discussion
We find it interesting that PXRDnet actually works very well, if not better, in cases where the information content was previously thought to be lower, e.g., 10 Å nanocrystal size. This provides an optimistic note for future computational crystallographic and materials science research, as it shows that problems that are harder for humans are not necessarily harder for data-driven and machine learning methods.
@@
The model is already out-performing in terms of structural complexity earlier attempts at nanoparticle structure solution [5, 47, 6] and more refinements to the approach are eminently possible.
@@ Methods
The files are made available as a dataset we call MP-20-PXRD which we include with our open source code. We hope that this will encourage the computational material science community to extend our work and explore the structure solution from PXRD problem even more.
We use MP-20-PXRD to train PXRDnet end-to-end. MP-20-PXRD contains 45,229 materials, and we use a 90-7.5-2.5% train-val-test split.
@@
We also use the 10 Å filter and Gaussian broadening on the experimental PXRD data, to make the patterns consistent with what was seen during train time by our model.
@@
To model the nano-scale finite size of the particles we assumed the diffraction pattern to be that of a long-range ordered crystal, convoluted with the Fourier transform of a window function on the PDF,
$\Pi(r/\tau)=\begin{cases}0&|r|>\frac{\tau}{2}\\ \frac{1}{2}&|r|=\frac{\tau}{2}\\ 1&|r|<\frac{\tau}{2}\\ \end{cases}$ (1)
where $r$ is the interatomic distance and $\tau$ is the nanomaterial width. We therefore convolute the diffraction pattern with its Fourier transform,
$\mathcal{F}\left\{\Pi\left(\frac{r}{\tau}\right)\right\}=\tau\frac{\text{sin}(\frac{Q\tau}{2})}{\frac{Q\tau}{2}}$ (2)
where $Q$ is the magnitude of the scattering vector [4]. The result of the convolution can be seen in Figure 1.
@@
We then simulate the PXRD pattern for each of the $100$ candidates with the PyMatGen package [38]. We simulate nanoscale shrinkage and apply Gaussian smoothing on these candidates’ simulated PXRD patterns, just as was done for the ground truth PXRD. Finally, the top $5$ candidates are selected based on the adherence (measured as $l_{1}$ loss) of simulated PXRD to the original, desired PXRD $y$. These $5$ candidates are treated as PXRDnet’s estimate of possible structure solutions of the molecule given the PXRD pattern and chemical composition. (In principle, we can have an arbitrarily high number of candidates, if we wish to expand our search.)
@@
Baselines
@@
To our knowledge, there exist no ab initio end-to-end structure solution methods for nanomaterials (i.e., PXRD pattern and formula as input, structure as output). Thus, to assess the ability of PXRDnet, we conduct two ablations.
Latent Space Search
@@
For our first ablation, we want to investigate how beneficial our latent code selection process is at creating materials that adhere to the given PXRD and formula information. Thus, we keep all the same parts of our generative structure solution process, except that we do not use PXRD or formula information to optimize the latent codes via backpropagation.
@@
Instead, we randomly sample the same number ($100$) of latent codes from the same distribution $N(0,I)$ as before. Then, we skip the latent code optimization, and go directly to the decoding process, which is conducted in exactly the same way as our method. That is, we use the same model with the same noise schedule, and we keep the teacher forcing, such that the atomic species and number of atoms are fixed to the true chemical formula. Finally, we do the same candidate filtering with PXRD simulation, to see which of the generated materials matches the ground truth PXRD pattern. Again, we keep the top $5$ candidates, and calculate evaluation metrics in exactly the same way as for our method.
Semi-Random Material Construction
For our next ablation, we want to investigate the importance of the atomic denoising process. So, in addition to removing latent code optimization, we also remove the atomic denoising process.
It is conducted as follows. We randomly sample $100$ latent vectors $z\sim N(0,I)$, the same number as in the regular material generation process. Without any optimization, we then use these latent vectors (via decoding from our trained MLP) to initialize the unit cells of the materials. As in the other evaluations, we assume that we know the chemical formula, and we randomly initialize the atomic positions, as in our method. The difference is that we do not conduct any denoising on these positions. Then, in exactly the same way as our method, we calculate the PXRD adherence between each of these $100$ materials and the ground truth PXRD, and choose the top $5$ candidates. From there, evaluation metrics are calculated in the same way as previously described.
This is semi-random, because while $z$ is randomly drawn from the latent distribution, once it is decoded by the MLP, it does contain priors about physically plausible unit cell shapes. For this reason, this is actually a generous baseline, because it makes use of a small portion of our trained model. Essentially, the only differences between this baseline and our method are (1) the lack of PXRD-guided latent code optimization; (2) the lack of denoising. All other parts are the same, making for a fair comparison.
Metrics
To quantify PXRDnet performance, we employ a set of rigorously defined metrics to compare candidate and base truth material similarity.
Match Rate
First, following prior work [28, 61, 62, 38, 22, 29, 23, 24, 25], we use StructureMatcher from PyMatGen with the following default tolerance settings: stol=0.3, angletol=5, ltol=0.2. We note that these tolerance settings are significantly stricter than those used to evaluate the original CDVAE model [28]. We report the rate at which StructureMatcher finds a match between the top PXRDnet candidate (as measured by simulated $l_{1}$ PXRD difference) and base truth material across $200$ materials sampled from the test set. Intuitively, the match rate is the percentage of PXRDnet predicted materials (where the prediction is the best item from the candidate set for a given PXRD pattern) that have approximately the same structure as the true material, within some small error thresholds.
PDF Correlation
We also report quantitative results on the similarity between simulated PXRD patterns for PXRDnet candidates and the target PXRD pattern. To quantify PXRD similarity, we convert the candidate and base truth $Q$-space PXRD patterns (without any sinc broadening or noise) to their pair-distribution-function counterparts via a sine Fourier transform [3]. Intuitively, pair distribution function measures the distribution of interatomic distances in a material. We then compute the Pearson correlation coefficient (PCC) [37] between the candidate and ground truth PDF patterns. A PCC of $1$ indidcates the curves are identical and a value of 0 that they curves are completely uncorrelated. PCC can go negative for anti-correlated curves, but this rarely happens in practice when comparing PDFs and so this measure de facto varies between 0 and 1 in these tests. This practice follows prior evaluation in the literature [32]. We report the best Pearson correlation value among the $5$ PXRDnet candidates.
R-Factor
To have a metric that operates directly on the reciprocal space diffraction patterns we also calculate the R-factor (residuals function) defined as,
@@
where $I_{\text{gt}}(Q)$ is the ground truth PXRD pattern in $Q$-space, and $I_{\text{pred}}(Q)$ is the PXRD pattern in $Q$-space simulated from the predicted crystalline structure. The PXRD patterns are computed on a grid of $8192$ evenly spaced samples in $Q$ space. Typically, $R_{wp}^{2}$ below $10\%$ can be considered to be a success [3].
@@
Finally, just as with the PDF correlation, we report the best $R_{wp}^{2}$ among the $5$ PXRDnet candidates for each material.
@@
Our intention was to randomly sample 20 materials for Rietveld refinement, which we did by uniformly selecting materials from the test set. For each of these materials, we picked a high-performing one of the top-5 candidates, which was then Rietveld refined.
@@
Since the lattice parameters and space group symmetry (all in P1) of the predictions are different than the ground truth, the simulated diffraction patterns of the predictions are typically different than those of the ground truth despite the topological similarity. We performed a constrained search over distortions of the predicted unit cell via refinements using the program TOPAS v7 [63]. Due to the different simulation engine, we first ensured a suitable profile to describe the peak-shape by fitting the ground truth model to the ground truth simulated data using a Gaussian convolution for the 100 Å and the 10 Å patterns. The resulting values were then fixed for refinements of PXRDnet candidates, which had been simulated with the same Gaussian filter as the ground-truth patterns. For each predicted structure, all lattice parameters were then allowed to refine within a window of +/- 20%  along with a scale factor, where the parameters were randomized after each convergence to escape local minima over 2000 iterations. Due to the low model symmetry and low angular range and high broadening of the patterns, we did not attempt to additionally refine the atom site positions. This is in contrast to the final refinement step in conventional structure solution from crystalline materials and so we do not expect to obtain $R_{wp}$ values as good as those from full refinements, even for correct structures.
@@
Finally, we compare the pre-refinement $R_{wp}^{2}$ to the post-refinement $R_{wp}^{2}$. A goal here is to look for correlations between metrics for structures that are successful pre and post refinement. If there is a strong correlation we can use pre-refinement metrics as a low cost proxy for models that will perform well post-refinement.
Although during the Rietveld refinement, we optimize $R_{wp}^{2}$ from the material degraded by concurrent sinc and Gaussian convolutions, for this comparison, we compare all the $R_{wp}^{2}$ (both pre- and post- refinement) without any simulated nanoscale broadening but with the low-sigma Gaussian broadening only. We feel that this gives a better assessment of the true quality of the structure solution. See "Metrics/R-Factor" for more details on the calculations.
@@ Acknowledgements
SJLB and HL proposed the research topic. GG and TLS proposed the machine learning methodology and wrote the code. GG, TLS, and SJLB conducted the theoretical nanomaterial shrinkage analysis. GG and SJLB procured the data. GG, SJLB, and HL designed the experiments. MWT conducted structural refinement. All authors wrote the manuscript and provided feedback.
@@
Work in the Lipson group was supported by U.S. National Science Foundation under AI Institute for Dynamical Systems grant 2112085. Work in the Billinge group was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences (DOE-BES) under contract No. DE-SC0024141.

### v2

@@ Results
Figure 5 shows $R_{wp}^{2}$ values for these candidates before and after the Rietveld refinement step, and Figure 4 shows the PXRD patterns pre- and post-refinement. The Rietveld refinement works very well for the 100 Å test cases, which have sharper Bragg peaks. In almost all of the cases (18 out of 20), the refinement brought $R_{wp}^{2}$ below 20%. In a majority of the cases (15 out of 20), the refinement brought $R_{wp}^{2}$ below 10%. There may still be some minor issues with the structure in these cases, but PXRDnet is clearly and consistently returning structures that are quite similar to the true one and likely close enough to yield a correct structure with some human input in each case.
It is also fascinating that the pre-refinement $R_{wp}^{2}$ of the candidates that led to the best post-refinement results were so high. We see some cases where a candidate with initial R-factor above $100\%$ got refined into near-perfect matches.
@@
The Rietveld refinements were less successful for the 10 Å nanoparticles, which have highly distorted diffraction patterns. While the refinement was usually helpful, it did not necessarily result in successful structures.
@@
This aspect of the work can be developed in the future, with better understanding of how to refine structural models in a more automated way, either against the PDF [2, 50] or against powder data, to obtain a higher structure solution rate.
Encouragingly, we find that PXRDnet overcomes the simulation-to-real gap, in that its success, as measured by both visual analysis (Figure 6) and quantitative metrics (Table 1), are comparable to those obtained from the simulated data.
[…]
We also note that the results displayed in Figure 6 are from the best (lowest $R_{wp}^{2}$) candidate post-Rietveld refinement, as measured by adherence to the experimentally observed PXRD – other candidates may not be as successful. Nonetheless, this is a fair representation because in real scenarios, this PXRD is available to compare against.
Figure 6: Experimental Data: Structure solutions from experimentally observed PXRD patterns in the IUCr database [51]. The leftmost column is the ground truth structure, the middle column is the predicted structure from PXRDnet (lowest $R_{wp}^{2}$ candidate post-Rietveld refinement), and the rightmost column is the comparison of the simulated PXRD (red) from TOPAS v7 [52] to the actual experimentally observed PXRD (green).
@@ Discussion
We find it interesting that PXRDnet actually works very well in cases where the information content was previously thought to be very low, e.g., 10 Å nanocrystal size. This provides an optimistic note for future computational crystallographic and materials science research, as it shows that problems that are harder for humans are not necessarily harder for data-driven and machine learning methods.
@@
From a machine learning perspective, we use an off-the-shelf CDVAE backbone [32] to demonstrate this proof of principle, sharing many design and architectural similarities with concurrent works [6, 7, 5] (although the others are not open-sourced). Thus, the novelty of our work is not in the machine learning advance, but rather in the fact that we chart a path to successfully solving the once-unsolved grand scientific challenge of structure determination of nanometer-sized materials.
The model is already out-performing in terms of structural complexity earlier attempts at nanoparticle structure solution [8, 40, 9], and more refinements to the approach are eminently possible. One example of this is that our model is based on a CDVAE [32] backbone, but there have been recent developments in graph diffusion models [26, 53, 54, 55] that could easily be swapped in for the CDVAE backbone.
Another future direction is to make the model robust to background signals (e.g., solvent, container). We fed our model experimental data from a colleague that had a significant signal from the sample container, which confused the model – performance improved once we subtracted the container signal.
@@ Methods
The files are made available as a dataset we call MP-20-PXRD which we include with our open source code. We hope that this will encourage the computational material science community to extend our work and explore the nanostructure solution from PXRD problem even more.
We use MP-20-PXRD to train PXRDnet end-to-end. MP-20-PXRD contains 45,229 materials, and we use a 90-7.5-2.5% train-validate-test split.
@@
We also use the 100 Å filter and Gaussian broadening on the experimental PXRD data, to make the patterns consistent with what was seen during train time by our model.
@@
To model the nano-scale finite size of the particles, we follow the approximation of a finite size particle as a periodic crystal windowed over a certain domain in real space. This does not capture non-space filling clusters, faceted morphologies, surface relaxations and so on, which we defer to future studies. We then make use of Fourier analysis [50] to derive the correct form for the diffraction pattern of a nano-scale crystal.
A physically principled and computationally feasible approximate model is to convolute the diffraction pattern with a sinc2 function [45]:
$I_{\text{nano}}=I_{\text{c}}\circledast\text{sinc}^{2}\left(\frac{Q\tau}{2}\right),$ (1)
where $Q$ is the magnitude of the scattering vector, $\tau$ is the size of the nanoparticle, $I_{c}$ is the diffraction pattern of the infinitely repeating crystal, and $I_{\text{nano}}$ is the diffraction of the nano-scale crystal. Intuitively, the sinc2 kernel has the desirable properties: (a) it is non-negative, thereby being physically plausible; (b) as $\tau$ decreases (corresponding to smaller crystalline size), the kernel gets wider, thereby broadening the peaks further. In particular, keeping in mind the Fourier relationship between the interatomic distance and the diffraction pattern, finite size effects can be seen as multiplying the distribution of interatomic distances with the window function
$\Pi(r/\tau)=\begin{cases}0&|r|>\frac{\tau}{2}\\ \frac{1}{2}&|r|=\frac{\tau}{2}\\ 1&|r|<\frac{\tau}{2}\\ \end{cases}.$ (2)
Then, going from real space to Fourier (i.e., diffraction pattern) space,
$\mathcal{F}\left[\Pi(r/\tau)\right]=\tau\cdot\text{sinc}\left(\frac{Q\tau}{2}\right).$ (3)
The leading $\tau$ factor can be ignored, since intensity is on an arbitrary scale.
Mathematically, this is only an approximation, but nonetheless a good one that appears in standard texts on the subject [45]. The result of the convolution can be seen in Figure 1.
Practical Considerations of Nanoscale Shrinkage
@@
We emphasize that the PXRD Regressor and Inverted Crystal Parameter Decoder in Figure 9 actually activate the same latent space encoding $z$. They can do this because of the combined loss in Equation 4: the latent code is passed through both forward models, but rather than updating the model parameters in the gradient descent procedure, we update the latent code values based on the loss. In this sense, inversion of the models is not in the feed-forward sense, it is in the sense that the loss function is used to update the latent code parameters, rather than the model parameters as would be traditionally done.
@@
We then simulate the PXRD pattern for each of the $100$ candidates with the PyMatGen package [56]. We simulate nanoscale shrinkage and apply Gaussian smoothing on these candidates’ simulated PXRD patterns, just as was done for the ground truth PXRD. Finally, the top $10$ candidates are selected based on the adherence (measured as $l_{1}$ loss) of simulated PXRD to the original, desired PXRD $y$. These $10$ candidates are treated as PXRDnet’s estimate of possible structure solutions of the molecule given the PXRD pattern and chemical composition. (In principle, we can have an arbitrarily high number of candidates, if we wish to expand our search.)
@@
Baseline: CDVAE Latent Space Search
@@
To our knowledge, there exist no ab initio end-to-end structure solution methods for nanomaterials that we could use as benchmarks (i.e., PXRD pattern and formula as input, structure as output). The existing structure solution works are: (a) only for non-nanosized materials and (b) do not publicly release their code [6, 7, 5]. Thus, to assess the ability of PXRDnet, we conduct an ablation on the main modification we made to the CDVAE model: the ability to condition on PXRD patterns.
@@
In particular, we want to investigate how beneficial our latent code selection process is at creating materials that adhere to the given PXRD and formula information. Thus, we keep all the same parts of our generative structure solution process, except that we do not use PXRD or formula information to optimize the latent codes via backpropagation.
@@
Instead, we randomly sample the same number ($100$) of latent codes from the same distribution $N(0,I)$ as before. Then, we skip the latent code optimization, and go directly to the decoding process, which is conducted in exactly the same way as our method. That is, we use the same model with the same noise schedule, and we keep the teacher forcing, such that the atomic species and number of atoms are fixed to the true chemical formula. Finally, we do the same candidate filtering with PXRD simulation, to see which of the generated materials matches the ground truth PXRD pattern. Again, we keep the top $10$ candidates, and calculate evaluation metrics in exactly the same way as for our method.
Metric: R-Factor
To quantify PXRDnet performance, we use the R-Factor (residuals function), which is the standard metric used by crystallographers when determining the success of their structure solutions [3, 4]. It operates directly on the reciprocal space diffraction patterns: the advantage of this is that when deployed in real-world experimental settings, it is directly observable.
We calculate the R-factor (residuals function) defined as,
@@
where $I_{\text{gt}}(Q)$ is the ground truth PXRD pattern in $Q$-space, and $I_{\text{pred}}(Q)$ is the PXRD pattern in $Q$-space simulated from the predicted crystalline structure. (Sometimes, peaks corresponding to certain Q-values are up-weighted in the calculation [4]; but for simplicity, we give every point equal weight.) The PXRD patterns are computed on a grid of $8192$ evenly spaced samples in $Q$ space. Typically, $R_{wp}^{2}$ below $10\%$ can be considered to be a success [3].
@@
We note that this way of computing the R-factor is an unusually strict metric since we score model success by comparing to data of higher resolution than was actually shown to the model.
Finally, we report the best $R_{wp}^{2}$ among the $10$ PXRDnet candidates for each material. This is reasonable, because as stated earlier, the ground truth PXRD can be directly observed in experimental settings.
@@
We uniformly selected $20$ materials for Rietveld refinement from the test set. For each of these materials, we took the ten best candidates generated by PXRDnet, and refined all of them.
@@
Since the lattice parameters and space group symmetry (all in P1) of the predictions are different than the ground truth, the simulated diffraction patterns of the predictions are typically different than those of the ground truth despite the topological similarity. We performed a constrained search over distortions of the predicted unit cell via refinements using the program TOPAS v7 [52]. Due to the different simulation engine, we first ensured a suitable profile to describe the peak-shape by fitting the ground truth model to the ground truth simulated data using a Gaussian convolution for the 100 Å and the 10 Å patterns. The resulting values were then fixed for refinements of PXRDnet candidates, which had been simulated with the same Gaussian filter as the ground-truth patterns. For each predicted structure, all lattice parameters were then allowed to refine within a window of +/- 20%  along with a scale factor, where the parameters were randomized after each convergence to escape local minima over 2000 iterations. Due to the low model symmetry and low angular range and high broadening of the patterns, we did not attempt to additionally refine the atom site positions. This is in contrast to the final refinement step in conventional structure solution from crystalline materials and so we do not expect to obtain $R_{wp}^{2}$ values as good as those from full refinements, even for correct structures. (We note that TOPAS uses a slightly different calculation of $R_{wp}^{2}$ that incorporates statistical weighting, but nonetheless, good results with the TOPAS calculation correspond to good results with our calculation.)
@@
Finally, we compare the pre-refinement $R_{wp}^{2}$ to the post-refinement $R_{wp}^{2}$. A goal here is to look for correlations between metrics for structures that are successful pre and post refinement. If there is a strong correlation we can use pre-refinement metrics as a low cost proxy upper bound for models that will perform well post-refinement.
Although during the Rietveld refinement, we optimize $R_{wp}^{2}$ from the material degraded by concurrent sinc and Gaussian convolutions, for this comparison, we compare all the $R_{wp}^{2}$ (both pre- and post- refinement) without any simulated nanoscale broadening but with the low-sigma Gaussian broadening only. We believe that this gives a better assessment of the true quality of the structure solution.
@@ Acknowledgements
SJLB and HL proposed the research topic. GG and TLS proposed the machine learning methodology and wrote the code. GG, TLS, MV, and SJLB conducted the nanomaterial shrinkage analysis. GG and SJLB procured the data. GG, SJLB, and HL designed the experiments. MWT conducted structural refinement. All authors wrote the manuscript and provided feedback.
@@
Work in the Lipson group was supported by U.S. National Science Foundation under AI Institute for Dynamical Systems grant 2112085. Work in the Billinge group was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences (DOE-BES) under contract No. DE-SC0024141. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0025528 to G. Guo.